OpenStudy (anonymous):

105in = ______ft ____in

OpenStudy (anonymous):

1 ft = 12 inches so 1 inch = 1/12 of a foot so to convert inches to foot we divide the inches by 12 105 in = 105/12 ft = 8 ft 9 in

OpenStudy (anonymous):

8 feet, 9 inches